# Highly efficient quantum Stirling engine using multilayer Graphene

### v1

@@ I Introduction
Quantum thermodynamics examines how the fundamental principles of energy transfer, heat exchange, and entropy apply to small systems where quantum effects such as coherence, entanglement, spin correlations, and discrete energy levels give rise to thermodynamical properties with no classical counterpart, and where quantum statistics provides a natural framework to describe these phenomena Vinjanampathy and Anders (2016); Kosloff (2013); Deffner and Campbell (2019); Millen and Xuereb (2016); Binder et al. (2015, 2018); Myers et al. (2022); Zanin et al. (2019); Vieira et al. (2023); De Chiara et al. (2018); Castorene et al. (2025a); Ford and O’Connell (2005); Liuzzo-Scorpo et al. (2016); Syros (1999); Brandao et al. (2015).
@@
In this context, considerable attention has been devoted to understanding how quantum systems, such as trapped ions, superconducting qubits, quantum dots, and low-dimensional electronic materials and their diverse behaviors can be harnessed to implement heat engines capable of producing useful work. Among quantum thermal machines, the Stirling cycle has attracted particular interest due to its potential to achieve high efficiency in regimes where thermalization and isochoric processes can be controlled with precision Saygin and Sisman (2001); Sisman and Saygin (2001); Cruz et al. (2023); Raja et al. (2021); Cakmak and Sedehi (2023); Castorene et al. (2024); Yin et al. (2018); Papadatos (2023); Wu et al. (1998); Rastegar-Sedehi et al. (2025); Huang et al. (2002); Xia et al. (2024); Li et al. (2025); Araya et al. (2023); Chatterjee et al. (2021); Zhang et al. (2024); Koyanagi and Tanimura (2024). In light of these considerations, theoretical and experimental investigations have shown that quantum heat engines based on discrete energy spectra may surpass their classical counterparts in specific regimes, particularly when the working substance exhibits a tunable energy level structure and distinctive statistical properties Peña et al. (2019); Deffner (2018). Within this framework, graphene and its multilayer stacks under perpendicular magnetic fields stand out as promising platforms due to their unique Landau level spectra Castro Neto et al. (2009); Goerbig (2011); McCann and Fal’ko (2006); Min and MacDonald (2008); Yuan et al. (2011), combined with their ease of fabrication and precise experimental control Mani et al. (2019); Singh and Benjamin (2021); Myers et al. (2023); Koshino and Ando (2008); Nakamura and Hirasawa (2008); Ghai et al. (2024); Riffo et al. (2024); Rickhaus et al. (2012); Jahani et al. (2023); Valenzuela et al. (2015); Peng (2011); Bao et al. (2011).
In this regard, several studies have proposed graphene-based systems, including quantum dots and multilayer structures, as candidates for quantum engines Josefsson et al. (2018); Josefsson and Leijnse (2020); Erdman et al. (2017); Josefsson et al. (2019); Arroyo-Gascon et al. (2023); Kalugin et al. (2017); Pelc et al. (2015); Suarez Morell and Foa Torres (2012); Esposito et al. (2010); De and Muralidharan (2016); Du et al. (2020). It is important to emphasize that many previous studies rely exclusively on Boltzmann statistics and employ simplified descriptions that neglect the fermionic nature of the electronic working fluid in graphene Peña et al. (2020); Myers et al. (2023, 2022); Singh and Benjamin (2021). In contrast, the present analysis explicitly incorporates the Fermi–Dirac statistics of charge carriers under realistic doping and experimental conditions. Moreover, the large degeneracy of the Landau levels implies that the number of participating electrons can be very high, providing a unique framework to explore quantum corrections to thermodynamic quantities arising from the discreteness of the Landau-level spectrum while remaining within the thermodynamic limit. As a result, thermal fluctuations are strongly suppressed, making it possible to isolate genuine quantum effects from those associated with finite-size fluctuations.
@@
This study examines the quantum Stirling cycle using monolayer, AB-stacked bilayer, and ABC-stacked trilayer graphene under perpendicular magnetic fields as the working medium. By incorporating Fermi–Dirac statistics into the thermodynamic description, the analysis provides a realistic account of electronic population dynamics, particularly at low temperatures and finite doping, extending beyond the idealized charge-neutrality point and enhancing the relevance of the results to experimental conditions. Within a grand-canonical formalism, thermodynamic quantities such as internal energy and entropy are evaluated, accounting for the symmetries specific to each multilayer configuration.
The chemical potential $\mu$ is determined self-consistently throughout the cycle to ensure particle-number conservation, enabling a consistent realization of the Stirling cycle with the magnetic field as the external control parameter and the system evolving through isothermal and isomagnetic strokes. The analysis reveals that the distinct Landau levels spectra of monolayer, bilayer, and trilayer graphene systems determine the heat exchanged during the cycle and, consequently, the resulting efficiency and performance. Among these systems, AB-stacked bilayer graphene provides the most favorable regime, combining high efficiency with finite work output.
@@ II.1 Landau Level Spectrum in Multilayer Graphene
At low electronic doping, within the low-energy approximation where a Dirac-like equation provides an accurate description of the electronic properties of graphene, the Landau level spectra of multilayer structures in a perpendicular magnetic field can be expressed through simple analytical formulas McCann and Fal’ko (2006); Min and MacDonald (2008); Yuan et al. (2011). These relations are written in terms of the Fermi velocity $v_{F}\simeq 3\times 10^{6}$ [m/s], the elementary charge $e$, the Planck constant $\hbar$, the Landau level index $n$, and the external magnetic-field strength $B$.
II.1.1 Landau Level Spectrum of Monolayer Graphene
Monolayer graphene is a single layer of $sp^{2}$-hybridized carbon atoms arranged in a two-dimensional honeycomb lattice. The primitive unit cell of this structure contains two inequivalent sites, A and B, corresponding to the filled and empty circles in Fig. 1(a). The corresponding reciprocal lattice is also honeycomb, featuring two inequivalent points, $K$ and $K^{\prime}$, in the Brillouin zone Castro Neto et al. (2009). Each carbon atom contributes one electron to the valence and conduction bands. Remarkably, the low-energy electronic structure exhibits a linear dispersion relation, $\varepsilon_{\bf k}=\hbar v_{F}|{\bf k}|$, centered at the $K$ and $K^{\prime}$ symmetry points. This behavior is characteristic of massless Dirac fermions Castro Neto et al. (2009).
When subjected to a perpendicular magnetic field $B$, these ultrarelativistic quasiparticles exhibit a unique Landau-level spectrum given by Castro Neto et al. (2009)
$\varepsilon^{(1)}_{n,\pm}=\pm\hbar\omega_{c}^{(1)}\sqrt{n},\quad n=0,1,2,\ldots,$ (1)
where the characteristic cyclotron frequency is
$\omega_{c}^{(1)}=\sqrt{\frac{2ev_{F}^{2}B}{\hbar}}.$ (2)
A distinctive feature of this spectrum is the presence of a zero-energy Landau level ($n=0$) shared equally by electrons and holes. Unlike conventional two-dimensional electron gases, here the Landau levels are not equally spaced but instead scale proportional to $\sqrt{nB}$, a direct consequence of the massless Dirac-like Hamiltonian that effectively describes the electronic dynamics at low energies and at $B=0$.
These features give rise to a fourfold-degenerate zero-energy state, originating from the spin and valley degrees of freedom, while higher Landau levels preserve the same degeneracy. This aspect will be further addressed in the following section, where it plays a central role in the formulation of the grand partition function. The resulting spectral asymmetry and nonuniform level spacing have profound implications for both the quantum Hall effect and the thermodynamic behavior of monolayer graphene Koshino and Ando (2007); Rickhaus et al. (2012); Jahani et al. (2023); Jellal (2016).
II.1.2 Landau Level Spectrum of AB-Stacked Bilayer Graphene
In AB-stacked bilayer graphene, shown in Fig. 1(b), the low-energy excitations correspond to massive chiral quasiparticles, whose pseudospin is locked to the momentum orientation due to the interlayer coupling. Under the application of a perpendicular magnetic field $B$, their corresponding Landau level spectrum is described by McCann and Fal’ko (2006)
$\varepsilon^{(2)}_{n,\pm}=\pm\hbar\omega_{c}^{(2)}\sqrt{n(n-1)},\quad n=0,1,2,\dots,$ (3)
where the effective cyclotron frequency $\omega_{c}^{(2)}$ and effective mass $m^{*}$ are, respectively
$\omega_{c}^{(2)}=\frac{eB}{m^{*}}\qquad\mbox{and}\qquad m^{*}=\frac{t_{\perp}}{2v_{F}^{2}}.$ (4)
Here, the dominant interlayer hopping parameter is $t_{\perp}\simeq 0.39\,\text{eV}$, which yields an effective mass $m^{*}\simeq 0.0343\,m_{e}$. In contrast to monolayer systems, not only $n=0$ forms a zero-energy Landau level, but also $n=1$, leading to a twofold degeneracy of the zero mode that is further enlarged by spin and valley symmetries. For $n\geq 2$, the Landau levels are unequally spaced and increase approximately linearly with $B$, in contrast to the $\sqrt{B}$-scaling observed in the monolayer. This spectral structure plays a central role in determining the thermodynamic and transport properties of bilayer graphene under perpendicular magnetic fields McCann and Fal’ko (2006).
II.1.3 Landau Level Spectrum of ABC-Stacked Trilayer Graphene
In ABC-stacked (rhombohedral) trilayer graphene, shown in Fig. 1(c), the low-energy bands follow a cubic dispersion $E\propto k^{3}$, in contrast to the linear and quadratic cases discussed previously. Under a perpendicular magnetic field $B$, the Landau levels are given by Bao et al. (2011)
$\varepsilon_{n,\pm}^{(3)}=\pm\hbar\omega_{c}^{(3)}\sqrt{n(n-1)(n-2)},\quad n\in\{0,1,2,\dots\},$ (5)
where the cyclotron frequency is
$\omega_{c}^{(3)}=\frac{\left(2eBv_{F}^{2}\right)^{3/2}\sqrt{\hbar}}{t_{\perp}^{2}}$ (6)
Here, $v_{F}$ denotes the Fermi velocity and $t_{\perp}\simeq 0.3\,\text{eV}$ is the dominant interlayer hopping parameter.
This spectrum exhibits the characteristic scaling $\varepsilon_{n}\propto B^{3/2}$, consistent with cubic-dispersion models reported in the literature. The first three levels ($n=0,1,2$) collapse to zero energy, increasing the zero-mode degeneracy. For $n\geq 3$, the levels become non-equidistant and increasingly spaced, while remaining compressed compared to bilayer systems due to the $\sqrt{n(n-1)(n-2)}$ dependence.
Since the Landau-level spectra are analytically well established for all three systems, a sufficiently large yet finite number of levels is included in the calculations to ensure numerical convergence over the relevant temperature and magnetic-field ranges.
@@ III.1 Grand canonical formalism
The thermodynamic properties of electrons in graphene multilayer systems under perpendicular magnetic fields are naturally described within the grand-canonical ensemble, where both the temperature $T$ and the chemical potential $\mu$ serve as control parameters. In this framework, the central quantities of interest are the internal energy $U$, the entropy $S$, and the particle number $\langle N\rangle$, which together determine the efficiency and work output of thermal cycles.
@@
For multilayer graphene with quantized Landau levels $\varepsilon_{n,\pm}^{(i)}$ (with $i=1,2,3$ denoting monolayer, bilayer, and trilayer), the grand-canonical partition function can be written explicitly as
$\mathcal{Z}^{(i)}=\prod_{n,\pm}\left[1+e^{-\beta\left(\varepsilon_{n,\pm}^{(i)}-\mu\right)}\right]^{g_{n}^{(i)}},$ (7)
For the Landau-level energies defined in Eqs. (1), (3), and (5), the corresponding degeneracies $g_{n}^{(i)}$ are given by
  • For all levels with $n\geq 1$ have degeneracy $g_{n}^{(i)}=4$.
  • The zero-energy level ($n=0$) exhibits enhanced degeneracy:
    $g_{0}^{(1)}=4,\qquad g_{0}^{(2)}=8,\qquad g_{0}^{(3)}=12.$ (8)
Taking these degeneracies into account, the grand potential can be expressed as
$\Omega^{(i)}=-k_{B}T\sum_{n,\pm}g_{n}^{(i)}\ln\left[1+e^{-\beta\left(\varepsilon_{n,\pm}^{(i)}-\mu\right)}\right],$ (9)
from which the thermodynamic quantities follow as
@@
Equivalently, one may start from the Fermi–Dirac distribution, which gives the occupation of each single-particle state,
$f(\varepsilon,\mu,T)=\frac{1}{e^{\beta(\varepsilon-\mu)}+1},\qquad\beta=\frac{1}{k_{B}T},$ (11)
together with the density of states (DOS),
$\nu^{(i)}(\varepsilon)=\sum_{n,\pm}g_{n}^{(i)}\,\delta(\varepsilon-\varepsilon_{n,\pm}^{(i)}).$ (12)
This allows the same observables to be written as
$\expectationvalue{N(\mu,T)}$ $=\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(i)}(\varepsilon)\,f(\varepsilon,\mu,T),$ (13)
$U(\mu,T)$ $=\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(i)}(\varepsilon)\,\varepsilon\,f(\varepsilon,\mu,T),$ (14)
$S(\mu,T)$ $=-k_{B}\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(i)}(\varepsilon)\,\left[f\ln f+(1-f)\ln(1-f)\right].$ (15)
Both approaches are formally equivalent: the discrete-level formulation based on the grand potential and the continuum representation using the DOS combined with the Fermi–Dirac distribution, both yielding to identical thermodynamic quantities. It is worth emphasizing that the DOS representation makes explicit the role of Landau-level degeneracies, particularly the zero-energy states in all systems. For the purposes of this work, we adopt the latter approach in our calculations.
III.2 Energy Scales and Parametrization of the Chemical Potential to Conserve Particle Number

### v2

@@ I Introduction
Quantum thermodynamics examines how the fundamental principles of energy transfer, heat exchange, and entropy apply to small systems where quantum effects such as coherence, entanglement, spin correlations, and discrete energy levels give rise to thermodynamical properties with no classical counterpart, and where quantum statistics provides a natural framework to describe these phenomena Vinjanampathy and Anders (2016); Kosloff (2013); Deffner and Campbell (2019); Millen and Xuereb (2016); Binder et al. (2015, 2018); Myers et al. (2022); Zanin et al. (2019); Vieira et al. (2023); De Chiara et al. (2018); Castorene et al. (2025a); Liuzzo-Scorpo et al. (2016); Brandao et al. (2015).
@@
In this context, considerable attention has been devoted to understanding how quantum systems, such as trapped ions, superconducting qubits, quantum dots Esposito et al. (2010); Erdman et al. (2017); Josefsson et al. (2018, 2019); Josefsson and Leijnse (2020); Du et al. (2020), and low-dimensional electronic materials and their diverse behaviors can be harnessed to implement heat engines capable of producing useful work. Among quantum thermal machines, the Stirling cycle has attracted particular interest due to its potential to achieve high efficiency in regimes where thermalization and isochoric processes can be controlled with precision Saygin and Sisman (2001); Sisman and Saygin (2001); Cruz et al. (2023); Raja et al. (2021); Cakmak and Sedehi (2023); Castorene et al. (2024); Yin et al. (2018); Papadatos (2023); Wu et al. (1998); Rastegar-Sedehi et al. (2025); Huang et al. (2002); Xia et al. (2024); Li et al. (2025); Araya et al. (2023); Chatterjee et al. (2021); Zhang et al. (2024); Koyanagi and Tanimura (2024). In light of these considerations, theoretical and experimental investigations have shown that quantum heat engines based on discrete energy spectra may surpass their classical counterparts in specific regimes, particularly when the working substance exhibits a tunable energy level structure and distinctive statistical properties Peña et al. (2019); Deffner (2018).
Graphene and its multilayer stacks, when subjected to perpendicular magnetic fields Castro Neto et al. (2009); McCann and Fal’ko (2006); Min and MacDonald (2008); Yuan et al. (2011); Goerbig (2011) stand out as promising platforms for quantum heat engines. Beyond their ease of fabrication and high degree of experimental control Koshino and Ando (2008); Nakamura and Hirasawa (2008); Ghai et al. (2024); Riffo et al. (2024); Rickhaus et al. (2012); Jahani et al. (2023); Valenzuela et al. (2015); Peng (2011); Bao et al. (2011), manifestations of their quantized spectrum are robust against temperature, since these systems exhibit quantum Hall plateaus even at room temperatures under sufficiently high magnetic fields Novoselov et al. (2007) – a critical feature for practical applications. Furthermore, the large degeneracy of the Landau levels involves a macroscopic number of participating electrons, $N$, even in micron-size flakes Goerbig (2011) This provides a unique framework to explore quantum corrections to thermodynamic quantities stemming from the discrete Landau-level spectrum, while remaining within the thermodynamic limit. Consequently, the small-$N$ thermal fluctuations typically found in few-body systems (both classical and quantum) are suppressed, making it possible to isolate genuine quantum effects. Very importantly, when driven by a varying magnetic field $B$, these quantum systems - unlike most others - avoid the dissipation associated to Landau-Zener transitions Kayanuma and Nakayama (1998); Dai et al. (2025), as Landau levels do not cross with changes in $B$. Finally, like other two-dimensional materials, the chemical potential (or doping) can be precisely tuned via electrostatic gating, offering a robust mechanism for controlling the engine’s working fluid.
Several studies have proposed graphene-based systems, including quantum dots and multilayer structures, as promising candidates for quantum engines Mani et al. (2019); Singh and Benjamin (2021); Myers et al. (2023); Arroyo-Gascon et al. (2023); Kalugin et al. (2017); Pelc et al. (2015); Suarez Morell and Foa Torres (2012); Soufy and Benjamin (2025). While these pioneering works offer valuable insights, they predominantly rely on Boltzmann statistics Peña et al. (2020); Myers et al. (2023, 2022); Singh and Benjamin (2021). By neglecting the fermionic nature of the electronic working fluid and assuming charge neutrality, the lack of quantitative accuracy of their results and the narrow focus of their setting call for a reexamination.
@@
This study examines a quantum Stirling cycle implemented in monolayer, AB-stacked bilayer, and ABC-stacked trilayer graphene under perpendicular magnetic fields as the working medium. Employing a low-energy Landau-level description and Fermi–Dirac statistics at finite doping, we correct the previous studies, expand they scope beyond the charge-neutrality point, and provide a thermodynamic analysis more closely aligned with experimentally accessible conditions Pan et al. (2016); Moore et al. (2025). Within a grand-canonical framework, we evaluate key observables such as internal energy and entropy while explicitly accounting for the degeneracies specific to each multilayer configuration.
A central motivation of this work is to clarify how the microscopic Landau–level structure of mono-, bi-, and trilayer graphene shapes the performance of a quantum Stirling cycle. We show that the stacking-dependent spectra of these multilayers can naturally produce parameter regimes in which the Stirling efficiency approaches the Carnot bound while still delivering a finite amount of energy per cycle in the reversible, quasistatic limit (and thus zero power), without relying on quantum criticality or ideal regenerative mechanisms. This highlights the role of stacking order as a physically meaningful variable for controlling the thermodynamic response of quantum heat engines.
@@ II.1 Landau Level Spectrum in Multilayer Graphene
At low electronic doping, the low-energy LL spectra of mono-, AB-bilayer, and ABC-trilayer graphene in a perpendicular magnetic field (as shown in Fig. 1) can be cast into a single compact form that highlights the multilayer scalings McCann and Fal’ko (2006); Min and MacDonald (2008); Yuan et al. (2011); Castro Neto et al. (2009); Bao et al. (2011).
Introducing a stacking index $J$ ($J=1$ monolayer, $J=2$ AB bilayer, $J=3$ ABC trilayer), we write
$\varepsilon^{(J)}_{n,\pm}=\pm\hbar\omega_{c}^{(J)}\sqrt{\prod_{j=0}^{J-1}(n-j)},\qquad n=0,1,2,\ldots,$ (1)
where $n$ is the LL index and $\pm$ indicate conduction and valence band LLs. The $J$-dependent cyclotron frequency $\omega_{c}^{(J)}$ encodes the microscopic constants of each system. Equation (1) immediately yields the characteristic magnetic-field scaling
$\varepsilon^{(J)}_{n}\propto B^{J/2},$ (2)
and predicts $J$ zero-energy modes ($n=0,1,\ldots,J-1$). Consequently, the effect spin, valley and zero-energy level degeneracies results in a combined degeneracy $g_{n}^{(J)}=4$ for all levels with $n\geq 1$ and
$g_{0}^{(1)}=4,\qquad g_{0}^{(2)}=8,\qquad g_{0}^{(3)}=12.$ (3)
for $n=0$.
For the three cases addressed here, the corresponding prefactors are:
$J=1:\quad\varepsilon^{(1)}_{n,\pm}$ $=\pm\hbar\omega_{c}^{(1)}\sqrt{n},$ (4)
$\omega_{c}^{(1)}$ $=\sqrt{\frac{2ev_{F}^{2}B}{\hbar}},$
$J=2:\quad\varepsilon^{(2)}_{n,\pm}$ $=\pm\hbar\omega_{c}^{(2)}\sqrt{n(n-1)},$ (5)
$\omega_{c}^{(2)}$ $=\frac{eB}{m^{*}},\qquad m^{*}=\frac{t_{\perp}}{2v_{F}^{2}},$
$J=3:\quad\varepsilon^{(3)}_{n,\pm}$ $=\pm\hbar\omega_{c}^{(3)}\sqrt{n(n-1)(n-2)},$ (6)
$\omega_{c}^{(3)}$ $=\frac{\left(2eBv_{F}^{2}\right)^{3/2}\sqrt{\hbar}}{t_{\perp}^{2}}.$
@@ III.1 Grand canonical formalism
In recent studies using graphene as a thermal machine, the thermodynamic behavior of graphene multilayers has been investigated through thermal states based on the canonical ensemble and Landau-level spectra Soufy and Benjamin (2025); Myers et al. (2023), neglecting the fermionic nature of the working fluid. In contrast, in this manuscript the thermodynamic properties of electrons in graphene multilayer systems under perpendicular magnetic fields are naturally described within the grand-canonical ensemble using Fermi–Dirac statistics, where both the temperature $T$ and the chemical potential $\mu$ serve as control parameters.
In this framework, the central quantities of interest are the internal energy $U$, the entropy $S$, and the particle number $\langle N\rangle$, which together determine the efficiency and work output of thermal cycles.
@@
For multilayer graphene with quantized Landau levels $\varepsilon_{n,\pm}^{(J)}$ (with $J=1,2,3$ denoting monolayer, bilayer, and trilayer), the grand-canonical partition function can be written explicitly as
$\mathcal{Z}^{(J)}=\prod_{n,\pm}\left[1+e^{-\beta\left(\varepsilon_{n,\pm}^{(J)}-\mu\right)}\right]^{g_{n}^{(J)}},$ (7)
Taking the degeneracies into account, the grand potential can be expressed as
$\Omega^{(J)}=-k_{B}T\sum_{n,\pm}g_{n}^{(J)}\ln\left[1+e^{-\beta\left(\varepsilon_{n,\pm}^{(J)}-\mu\right)}\right],$ (8)
where $\beta=\frac{1}{k_{B}T}$. The grand potential gives
@@
Equivalently, one may start from the Fermi–Dirac distribution, which gives the occupation of each single-particle state $f(\varepsilon,\mu,T)=[e^{\beta(\varepsilon-\mu)}+1]^{-1}$ together with the density of states (DOS),
$\nu^{(J)}(\varepsilon)=\sum_{n,\pm}g_{n}^{(J)}\,\delta(\varepsilon-\varepsilon_{n,\pm}^{(J)}),$ (10)
which does not yet include the LL orbital degeneracy, that is to be discussed in the next section. These allow the same observables to be written as
$\expectationvalue{N(T,B,\mu)}$ $=\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(J)}(\varepsilon)\,f(\varepsilon,\mu,T),$ (11)
$U(T,B,\mu)$ $=\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(J)}(\varepsilon)\,\varepsilon\,f(\varepsilon,\mu,T),$ (12)
$S(T,B,\mu)$ $=-k_{B}\int\limits_{-\infty}^{\infty}d\varepsilon\,\nu^{(J)}(\varepsilon)\,\left[f\ln f+(1-f)\ln(1-f)\right].$ (13)
Both approaches are formally equivalent: the discrete-level formulation based on the grand potential and the continuum representation using the DOS combined with the Fermi–Dirac distribution, both yielding to identical thermodynamic quantities. It is worth emphasizing that the DOS representation makes explicit the role of Landau-level degeneracies, particularly the zero-energy states in all systems. We adopt the latter approach for our calculations. To guarantee numerical stability across the entire parameter space, the sums over Landau levels are truncated at a cutoff index of $n=2000$, or when the relative difference between consecutive evaluations falls below $10^{-6}$. Crucially, for the specific temperature and magnetic field regimes considered, thermal population is confined to the lowest levels, with contributions from $n>20$ being exponentially suppressed. Consequently, the effective number of participating states is limited to $n\lesssim 20$, which is fully consistent with the $\delta$-function DOS representation employed in this work Vidarte and Lewenkopf (2022).
III.2 Energy Scales and Parametrization of the Chemical Potential